# Automated discovery and optimization of autonomous quantum error correction codes for a general open quantum system

### v2

@@ II Related Recent work
Other experiments include realization of a binomial code by Gertler et al. Gertler , a Gottesman-Kitaev-Preskill (GKP) code by Lachance-Quirion et al. LachanceQuirion , a qudit GKP code by Brock et al. Brock , all of which used 3-dimensional cavities coupled to superconducting qubits. A few years earlier, Leghtas et al. Leghtas had demonstrated state stabilization of two quantum states, although coherence between the two states was not preserved in that experiment. DeBry et al. DeBry and Li et al. Li2025 demonstrated AQEC in four-level codes in trapped ions.
@@
In the context of our work on the numerical discovery of AQEC codes, we should also mention the related work on the optimization of encoding, measurements and measurement-outcome-conditioned unitary operators in the standard measure-then-correct paradigm of QEC Reimpell ; Yamamoto2005 ; Fletcher2007 ; Yamamoto2007 ; Kosut ; Fletcher2008 ; PoulsenNautrup ; Bausch ; Olle ; Su ; Casanova . References Reimpell ; Yamamoto2005 ; Fletcher2007 ; Yamamoto2007 ; Kosut ; Fletcher2008 ; Olle addressed the question of whether various numerical optimization methods can efficiently find optimal QEC codes, including the code space and error correction operations. Poulsen Nautrup et al. PoulsenNautrup performed a numerical search that optimizes the connectivity graph in the surface code. Bausch et al. Bausch developed a machine-learning approach to achieve an optimal decoding strategy for noisy syndrome data when operating the surface code. References PoulsenNautrup ; Olle ; Su used reinforcement learning as the optimization tool for QEC code discovery. Casanova et al. Casanova use Riemannian optimization and included a performance metric that favours simple QEC codes over more complicated ones.
@@ III.1 Lindblad equation of motion
where $H$ is the Hamiltonian, which is responsible for unitary evolution, and $a_{j}$ are the jump operators of the different decay channels. The Lindblad equation is often written with decay rates appearing in the decay term. However, we choose the form in Eq. (1), in which the decay rates are absorbed into the jump operators, because this convention will be more convenient when we introduce the control and induced-decay terms. Specifically, since we will search for optimal values of the rates and jump operators, it is natural to combine them, rather than spend optimization resources on redundant variables.
@@ III.2 Defining target operation for optimal control calculation
This formula can be understood as follows: In a basis that contains the states $\left\{\left|\tilde{0}\right\rangle,\left|\tilde{1}\right\rangle\right\}$, each term inside the trace focuses on one of the four matrix elements of the density matrix in the code space and calculates how this term is degraded by the open-system dynamics described by $\tilde{L}$. The factor $1/4$ is used because the projector $\hat{P}$ has four matrix elements equal to one, with zeros everywhere else. The fidelity is equal to 1 if and only if the density matrix remains unaffected by the dynamical evolution, i.e. it experiences decoherence-free evolution. It should be noted that Eq. (6) treats the four matrix elements as if they were independent of each other, even though they are not. We find this formulation convenient when treating the density matrix as a one-dimensional vector, without worrying about the fact that this vector must obey the standard conditions for a physical density matrix. It should also be noted that there is no unique definition of the fidelity for the purpose of searching for AQEC codes. In principle, any function that is equal to a certain value when the quantum state is preserved but decreases when we move away from perfect state protection will be an acceptable function in the calculations. We find the fidelity in Eq. (6) to be a computationally convenient choice, as it is obtained using standard matrix operations.
@@ III.3 Temporal evolution of the fidelity
where $\beta_{1}$ and $\beta_{2}$ are chosen to give fidelities that are slightly below the initial value, such that the numerator gives a good approximation for the slope of the fidelity at early times. For realistic AQEC parameters, $\kappa$ should be largely insensitive to the exact values of $\beta_{1}$ and $\beta_{2}$, in addition to being insensitive to the fact that the fidelity starts at a value below $F=1$ even at $t$ values that are much smaller than the natural decay times.
@@ IV.2 Ququart with uniform decay rates
Figure 2: Progression of AQEC code with optimization iteration number $k$ for 4-level system with uniform decay rates. The infidelity (1F)(1-F), plotted in Panel (a), decreases steadily and approaches zero as a function of $k$, indicating that the algorithm is successful in finding a good AQEC code. Panel (b) shows that the code space approaches $\left\{\left|1\right\rangle,\left|3\right\rangle\right\}$. In Panel (c) the legend specifies only the highest two lines. In this specific run, by coincidence, the real parts of $b_{1,0}$ and $b_{3,2}$ did not grow significantly up to $k=10^{5}$. In Panel (d), the inset shows the real (cyan) and imaginary (gray) parts of $O_{1,3}$. This matrix element is expected to converge to zero in the optimal code. Most other matrix elements in $\hat{O}$ are expected to have a small effect on the fidelity, as explained in Appendix A. All plots in this figure indicate that the algorithm is progressing towards the 13 code but that it has not converged yet.
@@ IV.4 Photon loss in higher-dimensional qudit
where $\delta\theta=\pi/2-2\arctan\left(e^{-\gamma\delta t}\right)=\gamma\delta t+O(\delta t^{2})$ and $\hat{\sigma}_{y}^{(04)}=i\left(\left|4\right\rangle\left\langle 0\right|-\left|0\right\rangle\left\langle 4\right|\right)$. This operation can in turn be implemented by adding to the Hamiltonian the time-independent operator $\hat{O}=\gamma\hat{\sigma}_{y}^{(04)}$. In other words, $O_{0,4}=-O_{4,0}=-i$, with all other matrix elements in $\hat{O}$ equal to zero. It should be noted that the operator $\hat{\sigma}_{y}^{(04)}$ involves a four-photon process and therefore seems challenging to realize experimentally. There are, however, theoretical proposals and initial experimental results on implementing such multi-photon operations Kwon2022 ; Chang ; Eriksson .
@@ Appendix A Optimizing individual components of AQEC code
In this Appendix, we present additional results that elucidate the the performance of the optimization algorithm in optimizing the different components of the AQEC code. We use the same 4-level model as in Sec. IV.B. We perform a few different optimization tasks, focusing on one component at a time.
@@ Appendix B General qudit with uniform decay rates
We ran the optimization algorithm on the case of a 6-level system starting from a random seed, i.e. random guess for the AQEC code. In the majority of the runs, the fidelity reached $F\approx 0.9992$ after $10^{5}$ iterations. The maximum value reached in one run (out of ten runs) was $F=0.99938$. These fidelity values are much higher than the free-evolution value $F_{0}$ given in Sec. IV.A, but short of the value 0.999999 obtained with intuitive reasoning above. The optimization algorithm is clearly able to find good AQEC protocols. We expect that it will converge to the optimal code if it is run longer. One caveat here is that there are a large number of extremely good AQEC codes in this 6-level system, such that the algorithm can already reach fidelities at the level of several nines with codes that are drastically different from the possibly optimal one described above. Furthermore, convergence tends to be slow when the fidelity approaches the maximum, mainly because of the slow convergence of the induced decay matrix $\hat{b}$. An additional complication that does not affect this particular example but can be important for other decay models is the fact that higher levels generally decay faster and result in higher admixtures of error states. Even if these admixtures do not affect the effective decay rate, they contribute a small reduction to the fidelity, making it difficult for the algorithm to distinguish between many different codes that all have extremely high fidelities.

### v3

@@ II Related Recent work
Other experiments include the realization of a binomial code by Gertler et al. Gertler , a Gottesman-Kitaev-Preskill (GKP) code by Lachance-Quirion et al. LachanceQuirion , and a qudit GKP code by Brock et al. Brock , all of which used 3-dimensional cavities coupled to superconducting qubits. A few years earlier, Leghtas et al. Leghtas had demonstrated state stabilization of two quantum states, although coherence between the two states was not preserved in that experiment. DeBry et al. DeBry and Li et al. Li2025 demonstrated AQEC in four-level codes in trapped ions.
@@
In the context of our work on the numerical discovery of AQEC codes, we should also mention the related work on the optimization of encoding, measurements and measurement-outcome-conditioned unitary operators in the standard measure-then-correct paradigm of QEC Reimpell ; Yamamoto2005 ; Fletcher2007 ; Yamamoto2007 ; Kosut ; Fletcher2008 ; PoulsenNautrup ; Bausch ; Olle ; Su ; Casanova . References Reimpell ; Yamamoto2005 ; Fletcher2007 ; Yamamoto2007 ; Kosut ; Fletcher2008 ; Olle addressed the question of whether various numerical optimization methods can efficiently find optimal QEC codes, including the code space and error correction operations. Poulsen Nautrup et al. PoulsenNautrup performed a numerical search that optimizes the connectivity graph in the surface code. Bausch et al. Bausch developed a machine-learning approach to achieve an optimal decoding strategy for noisy syndrome data when operating the surface code. References PoulsenNautrup ; Olle ; Su used reinforcement learning as the optimization tool for QEC code discovery. Casanova et al. Casanova used Riemannian optimization and included a performance metric that favours simple QEC codes over more complicated ones.
@@ III.1 Lindblad equation of motion
where $H$ is the Hamiltonian, which is responsible for unitary evolution, and $\hat{a}_{j}$ are the jump operators of the different decay channels. The Lindblad equation is often written with decay rates appearing in the decay term. However, we choose the form in Eq. (1), in which the decay rates are absorbed into the jump operators, because this convention will be more convenient when we introduce the control and induced-decay terms. Specifically, since we will search for optimal values of the rates and jump operators, it is natural to combine them, rather than spend optimization resources on redundant variables.
@@ III.2 Defining target operation for optimal control calculation
This formula can be understood as follows: In a basis that contains the states $\left\{\left|\tilde{0}\right\rangle,\left|\tilde{1}\right\rangle\right\}$, each one of the four nonzero terms inside the trace focuses on one of the four matrix elements of the density matrix in the code space and calculates how this term is degraded by the open-system dynamics described by $\tilde{L}$. The factor $1/4$ is used because the projector $\hat{P}$ has four matrix elements equal to one, with zeros everywhere else. The fidelity is equal to 1 if and only if the density matrix remains unaffected by the dynamical evolution, i.e. it experiences decoherence-free evolution. It should be noted that Eq. (6) treats the four matrix elements as if they were independent of each other, even though they are not. We find this formulation convenient when treating the density matrix as a one-dimensional vector, without worrying about the fact that this vector must obey the standard conditions for a physical density matrix. It should also be noted that there is no unique definition of the fidelity for the purpose of searching for AQEC codes. In principle, any function that is equal to a certain value when the quantum state is preserved but decreases when we move away from perfect state protection will be an acceptable function in the calculations. We find the fidelity in Eq. (6) to be a computationally convenient choice, as it is obtained using standard matrix operations.
@@ III.3 Temporal evolution of the fidelity
where $\beta_{1}$ and $\beta_{2}$ are chosen to give fidelity values that are slightly below the initial value, such that the numerator gives a good approximation for the slope of the fidelity at early times. For realistic AQEC parameters, $\kappa$ should be largely insensitive to the exact values of $\beta_{1}$ and $\beta_{2}$, in addition to being insensitive to the fact that the fidelity starts at a value below $F=1$ even at $t$ values that are much smaller than the natural decay times.
@@ IV.2 Ququart with uniform decay rates
Figure 2: Progression of AQEC code with optimization iteration number $k$ for 4-level system with uniform decay rates. The infidelity (1F)(1-F), plotted in Panel (a), decreases steadily and approaches zero as a function of $k$, indicating that the algorithm is successful in finding a good AQEC code. The somewhat irregular behaviour of the slope suggests that the fidelity landscape has saddle-point-like features. Panel (b) shows that the code space approaches $\left\{\left|1\right\rangle,\left|3\right\rangle\right\}$. In Panel (c) the legend specifies only the highest two lines. In this specific run, by coincidence, the real parts of $b_{1,0}$ and $b_{3,2}$ did not grow significantly up to $k=10^{5}$. In Panel (d), the inset shows the real (cyan) and imaginary (gray) parts of $O_{1,3}$. This matrix element is expected to converge to zero in the optimal code. Most other matrix elements in $\hat{O}$ are expected to have a small effect on the fidelity, as explained in Appendix A. All plots in this figure indicate that the algorithm is progressing towards the 13 code but that it has not converged yet.
@@ IV.4 Photon loss in higher-dimensional qudit
where $\delta\theta=\pi/2-2\arctan\left(e^{-\gamma\delta t}\right)=\gamma\delta t+O(\delta t^{2})$ and $\hat{\sigma}_{y}^{(04)}=i\left(\left|4\right\rangle\left\langle 0\right|-\left|0\right\rangle\left\langle 4\right|\right)$. This operation can in turn be implemented by adding to the Hamiltonian the time-independent operator $\hat{O}=\gamma\hat{\sigma}_{y}^{(04)}$. In other words, $O_{0,4}=-O_{4,0}=-i$, with all other matrix elements in $\hat{O}$ equal to zero. It should be noted that the operator $\hat{\sigma}_{y}^{(04)}$ involves a four-photon process and is therefore challenging to realize experimentally. There are, however, theoretical proposals and initial experimental results on implementing such multi-photon operations, especially in superconducting circuits Kwon2022 ; Chang ; Eriksson .
@@ Appendix A Optimizing individual components of AQEC code
In this Appendix, we present additional results that elucidate the performance of the optimization algorithm in optimizing the different components of the AQEC code. We use the same 4-level model as in Sec. IV.B. We perform a few different optimization tasks, focusing on one component at a time.
@@ Appendix B General qudit with uniform decay rates
We ran the optimization algorithm on the case of a 6-level system starting from a random seed, i.e. random guess for the AQEC code. In the majority of the runs, the fidelity reached $F\approx 0.9992$ after $10^{5}$ iterations. The maximum value reached in one run (out of ten runs) was $F=0.99938$. These fidelity values are much higher than the free-evolution value $F_{0}$ given in Sec. IV.A, but short of the value 0.999999 obtained with intuitive reasoning above. The optimization algorithm is clearly able to find good AQEC protocols. We expect that it will converge to the optimal code if it is run longer. One caveat here is that there are a large number of extremely good AQEC codes in this 6-level system, such that the algorithm can already reach fidelity values at the level of several nines with codes that are drastically different from the possibly optimal one described above. Furthermore, convergence tends to be slow when the fidelity approaches the maximum, mainly because of the slow convergence of the induced decay matrix $\hat{b}$. An additional complication that does not affect this particular example but can be important for other decay models is the fact that higher levels generally decay faster and result in higher admixtures of error states. Even if these admixtures do not affect the effective decay rate, they contribute a small reduction to the fidelity, making it difficult for the algorithm to distinguish between many different codes that all have extremely high fidelities.